OpenStudy (anonymous):

Which is an equation in point-slope form of the line that passes through the points (7,5) and (-4,-1) ?

OpenStudy (aravindg):

First find slope.

OpenStudy (anonymous):

i think the slope is 5

OpenStudy (anonymous):

the equation is y=mx+b

OpenStudy (anonymous):

hello???

OpenStudy (aravindg):

How did you get slope as 5? :O

OpenStudy (anonymous):

ya!! isnt the slpoe 6/11

OpenStudy (anonymous):

using formula slope (m)= (y2-y1)/(x2-x1) m = (5-(-1))/(7-(-4)) = (5+1)/(7+4) = 6/11